OpenStudy (anonymous):

10=-2(s+2)+3(s-10) solve

OpenStudy (isaiah.feynman):

.

OpenStudy (anonymous):

Can anyone help me???

OpenStudy (anonymous):

10=-2(s+2)+3(s-10) First distribute 10=-2s-4+3s-30 combine like terms 10=s-34 add 34 to both sides 44=s

OpenStudy (anonymous):

Thx